Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard from; ax^2+bx+c=0,
Solution: 5, only solution.

Thanks!

Answer by rapaljer(4671)

Starting with x=5, this means that x-5=0

You can also write
(x-5)(x-5)=0
x^2-10x+25=0

Dr. Robert J. Rapalje, Retired